OpenStudy (anonymous):

Find an equation of the tangent to the curve at the point corresponding to the given value of the parameter . . .

OpenStudy (anonymous):

\[x = t^4 +1\] \[y = t^3 + t\] \[t = -1\]

OpenStudy (beginnersmind):

Can you solve this if you're given a function y(x)? Remember that y'(x)=dy/dx=(dy/dt)/(dx/dt)

OpenStudy (anonymous):

I can find the slope of the tangent, but what do I do with that information?

OpenStudy (anonymous):

OH WAIT. y - y0 = m(x - x0), right?

OpenStudy (beginnersmind):

yeah. Can you get y0 and x0?

OpenStudy (anonymous):

Yes! I just plug in t = -1?

OpenStudy (anonymous):

Awesome, thanks :)

OpenStudy (beginnersmind):

cool. What did you get?